BI preparation for LS2
Minutes from meeting of 4/9/18
PS complex TS2 (William)
Christos - 2 additional inspections, but this is an IMPACT from another group.
Stephane - would like to replace a camera in TT2 - no impact created yet, will do ASAP.
PS complex for LS2 (William)
Missing ECR for the BTP line - in the pipeline
43 magnets to be renovated in the PS during
15/1/19 to 10/8/2020 PS general access.
Schedules still not available
Need a list of LS2 activities.
SPS (Morad)
TS2 interventions - 4 impacts. no other requests
No new dates for SPS planning since ~May
Will start planning meetings
ECR for MOPOS - Manfred will check if this is approved.
Changes in LSS5 - Who is managing this?
TPSC4 BTV needs to get added to the list
Re-cabling in TT20 and motors in the BTV to be added. Who does the NA LS2 coordination?? - Morad will ask. In theory BTV, SEM and BLM are critical activities that will be accepted.
LHC-TS2 - Gerhard
TS2 - Mail sent with the dates. BLM exchange in pt7 on 20/9.
Generally access monday to friday.
LSS4 fluorecncec camera, BBQ schottky,
Timepix in point 1 - not sure about IMPACT, GS and JS will check.
Diamond BLM is now working, so perhaps not remove.
Visit for DISMAC - not a BI IMPACT
Manfred - may change activities from point 4 to point 2. There is an IMPACT already. This should be documented in some way.
LHC LS2 (Gerhard)
Warm-up is spaced - will have some access in point 4 before warm-up.
‘Easy' access until end of July 2020.
Christos notes that maintenance of elevators will block access in some areas.
BSRT need 10 days of daytime access in 2021 instrument check-out in Jan-Feb, with most needed in point 4. Preferably without too much overlap.
Change of cabling and boxes - reduced scope.
An additional BLM in point 7 will be added to the list.
BPM cry-cable inspection. We will ask to open the insulation vacuum sleeve, Manfred will give the positions to Gerhard who will make the request.
BSRT removal of redundant supports and cleaning of the vacuum chamber. Gerhard will organise a meeting to plan this activity.
PLAN 11374 for preliminary bSRT activities - Ray will CANCEL this in PLAN.
General discussion on work in progress
SPS BSRT - we may need to add this if we get money
SPS DCBCT - will need a new chamber when project is official.
Subjects for next meeting:
Schedules from planning team
Need more detail on when we have access and not.
List of BI activities in PLAN for the PS complex
Storage of equipment during dismantling
LSS5 re-re-organisation
Reports from LS2C meeting
Electrical lockout procedures
Dose planning - only the PS main unit renovation is being tracked by William
BCTW will move from point 5 to point 3. A separate ECR will be made. Ask Morad to put Tom in touch with ACE designers. Will also need a new support design..